OpenStudy (anonymous):

please can anybody help me with this thanks i suck at this type thanks starting from 100 ft away. your friend skates toward you and then passes you. she skates at a constant speed of 20ft/s. her distance d from you in feet after t seconds is given by d=/100-20t/. at what times is she 60 ft from you?

OpenStudy (anonymous):

60= 100 -20t -100 -100 -40/-20= -20t/-20 2=t

jimthompson5910 (jim_thompson5910):

Plug in d = 60 and solve for t. d=|100-20t| 60=|100-20t| |100-20t| = 60 100-20t = 60 or 100-20t = -60 -20t = 60-100 or -20t = -60-100 -20t = -40 or -20t = -160 t = -40/(-20) or t = -160/(-20) t = 2 or t = 8 So she is 60 ft away from you at 2 seconds and at 8 seconds

OpenStudy (anonymous):

Oh, didn't see absolute value ha :P